OpenStudy (scott208):

I enter the mall with 75 dollars. I spend $a dollars at American Eagle, which is $4 more than the amount I spent at Banana Republic. I also spent $15 less at American Eagle than I spent at Cole Haan. How much money do I have left after shopping at these three stores? (Answer in terms of a.) please help will fan and medal

OpenStudy (welshfella):

$ a at American Eagle $ a-4 spent at Banana $ a+15 at Cole Haan add these values up then subtract them from 75.

OpenStudy (scott208):

a+a-4+a+15=75

OpenStudy (scott208):

which is equal to 3a+11=75 correct?

OpenStudy (welshfella):

No that is incorrect She needs the amount left after spending the money.

OpenStudy (welshfella):

75 - a -(a - 4) - (a + 15) simplify that to get your answer.

OpenStudy (scott208):

ohh ok

OpenStudy (scott208):

86-3a?

OpenStudy (welshfella):

check the 86 don't forget to change the signs when you distribute negatives over the parentheses.

OpenStudy (welshfella):

for example - ( x - 4) = -x + 4

OpenStudy (welshfella):

- * - = +

OpenStudy (scott208):

75-a-a+4-a-15? and then simplify

OpenStudy (scott208):

64-3a?

OpenStudy (welshfella):

yes correct

OpenStudy (scott208):

so thats it?

OpenStudy (scott208):

thank you !

OpenStudy (welshfella):

thats the answer yw